need the final answers (without explanation) I'll rate you immediately

f Cooer oe tor the fist for monthe of the coming year Th f 90,0005 7000 Onavorage,50% ..fcredit sales are paid for in the month of the sale. 30% in the month ollne ing sale and the remainder is paid two months ater the month of the sale Assuming there aro ne had deht, the expected cash innow in March is: B) c) $122,000 $119.000 desired for how many S1O8,0 7 It there were 30,000 pounds of raw material on hand on January I, also 60.0 inventory at January 31, and 180,000 pounds pounds of raw material should be purchased in January are required for January production 150,000 pounds 120.000 pounds b240.000 pounds d 210,000 pounds 8, Ifthe activity level increases 10%, total variable costs will a remain the same c. decrease by less than 10%. b, d. increase 10% increase by more than 10%. 9. Arthur Company has a margin of safety percentage of 25% The break-even point is S300.000 and the varabl expenses are 45% of sales. Given this information, the net operating income is: A) $75,000 C) $15.000 B) S5 5,000. D) 341.250 10. Dock Company makes two products from a common input. Joint processing costs up to the split- point total $33,600 a year. The company allocates these costs to the joint products on the basis of the total sales values at the split-off point. Each product may be sold at the split-off point or processed further. Data concerning these products appear below Allocated joint processing costs Sales value at split-off point Costs of further processing Sales value after further processing Product X Product Y Total $16,800 $16,800 $33.600 $24,000 $24.000 $48,000 $15,000 $18,700 $33,700 S35,500 $45,100 $80.600 What is the net monetary advantage (disadvantage) of processing Product Y beyond the split-off A) $9,600 B) $2,400 C) $33,600 D) $26,400
Answer #1
6
March
Cash Sales 18000
Credit sales collection:
January sales 20000 =100000*20%
February sales 36000 =120000*30%
March sales 45000 =90000*50%
Total cash inflow 119000
Option C is correct
7
Pounds required for production 180000
Add: Desired ending inventory 60000
Less: Beginning inventory -30000
Pounds of raw materials to be purchased 210000
Option D is correct
8
Total variable costs will increase by 10%
Option D is correct
9
Sales over break even point 100000 =300000/75%*25%
Less: Variable expenses 45000 =100000*45%
Net operating income 55000
Option B is correct
10
Product Y
Sales value after further processing 45100
Less: Sales value at split-off point 24000
Incremental revenue 21100
Less: Costs of further processing 18700
Monetary advantage(disadvantage) 2400
Option B is correct